OpenStudy (anonymous):

Find the absolute max and min values of f(x)=(x^2)(e^x)....I know f'(x) is e^x(x^2+2x)...just can't get the critical points

OpenStudy (zarkon):

set equal to zero and solve...what is the problem?

OpenStudy (anonymous):

it said its 0 and e...I got 0 but not e

OpenStudy (zarkon):

x=0,-2

myininaya (myininaya):

\[e^x \neq 0\]

myininaya (myininaya):

so all you have to do is look at x^2+2x=0 and thats what zarkon did

OpenStudy (anonymous):

thats what I got but according to the answer e and 0 are the max & min

myininaya (myininaya):

e is crazy talk

OpenStudy (anonymous):

it was on my exam and I got it wrong and that was the correct answer

myininaya (myininaya):

is there an interval?

OpenStudy (anonymous):

[-1,1]

OpenStudy (zarkon):

look at the graph http://www.wolframalpha.com/input/?i=plot [x^2*e^x%2C+{x%2C-3%2C1}]

OpenStudy (zarkon):

AH....IS THE THE RESTRICTED DOMAIN?

OpenStudy (zarkon):

sorry for the caps ;)

OpenStudy (zarkon):

on [-1,1] e is the max value of the function

myininaya (myininaya):

so the critical numbers are 0 and -2 we don't care about -2 its not the interval [-1,1] \[f(-1)=(-1)^2e^{-1}=\frac{1}{e} ; f(1)=(1)^2e^{1}=e; f(0)=0\]

myininaya (myininaya):

largest is e so e is the abs max value 0 is smallest so 0 is the abs min value

OpenStudy (anonymous):

ooooo ok I get it

OpenStudy (zarkon):

you probably should have included the [-1,1] in the original problem ;)

myininaya (myininaya):

yep yep thats important info

OpenStudy (anonymous):

sorry